# A deconstruction of methods to derive one-point lensing statistics

### v1

@@ 2.5 Hydrodynamical simulations
Figure 1: Convergence PDFs for the different lensing methods considered in this study, for an angular resolution of 1.76 arcsec.
@@ 3 Results
We now present our results for the PDFs of convergence, shear, magnification, and magnitude. In the ray-tracing method employed, rays are calculated from the source to the observer, resulting in statistics computed in the image plane. In most cases, we subsequently translate these PDFs to the source plane, which is relevant to observations. Hereafter, the PDFs in the image and source planes are denoted by the subscripts I and S, respectively.
@@ 3.1 Lensing PDFs
Figure 2: Comparison of $\kappa$, $\gamma$ and $\mu$ PDFs (in the source plane) for different redshifts for the $N$-body and PINOCCHIO case for the angular resolution of 1.76 arcsec. $\texttt{\small PINOCCHIO}_{\textrm{halos}}$ (not shown here for compactness) exhibit similar behavior to PINOCCHIO. The shaded regions mark where baryonic corrections exceed 50% for $z=1$ (blue) or for $z=3$ (green). As can be seen the two methods agree reasonably well in the range where baryonic effects are under control. See also Fig. 3.
@@
Figure 3: Relative residual differences between approximate methods and the full $N$-body results in the source plane, for $z=1$ (solid lines) and $z=3$ (dashed lines). The blue (green) shaded regions represent the interval in which the baryonic correction is over 50% for $z=1$ ($z=3$).
@@
However, differences in the tails of the distributions are obscured by the linear scale used in Fig. 1. Therefore, in Fig. 2, we present a general comparison between the results from $N$-body and PINOCCHIO for $\kappa$, $\gamma$, and $\mu$ on a logarithmic scale, depicting four different redshifts for each method. The turboGL code is excluded from this comparison as it does not compute the shear PDFs so that $\mu$ can be inferred only in the weak-lensing regime. As previously discussed, our analysis focuses on the region where baryonic effects are maintained below the 50% level. The regimes where this is not the case at $z=1$ and $z=3$ are highlighted with a blue and green shade, respectively. It is evident that as the source distance increases, the convergence, shear, and magnification PDFs become wider (indicating higher variance), and the peaks of their distributions shift (indicating skewness). We will further explore these trends for convergence in Section 3.3, noting the general trend that lensing intensifies at higher redshifts due to the increased number of lens encounters. We find that both PINOCCHIO and $N$-body simulations are consistent with the theoretical model of randomly distributed point-mass lenses (Schneider et al., 1992), which predicts that when $\gamma<0.01$, the PDF is proportional to $\gamma^{2}$, and for $\mu\gg 1$, the PDF is proportional to $\mu^{-1}$. However, while the asymptotic behavior prediction of the randomly distributed points model for the $\gamma-$PDF happens within the range where the baryonic effects are subdominant, the prediction for the $\mu-$PDF occurs in the regime deeply dominated by baryonic physics and is not shown in Fig. 2. The behavior of the magnification PDF at this regime will be discussed in Sect. 3.2.
Figure 3 illustrates the differences between the PDFs from PINOCCHIO and turboGL compared to the $N$-body results. Both approximate methods exhibit a small excess of probability density at the peak of the distribution and a deficiency in the tail. These discrepancies stem from the limitations of the approximate methods in accurately reproducing the non-linear density field, likely due to an inadequate representation of substructures and halo concentrations which are prevalent in this non-linear regime and responsible for high-convergence events. We find that PINOCCHIO provides a better approximation than the stochastic method at $z=1$, but both methods perform comparably at $z=3$. However, both methods gradually fail for larger $\kappa$, exhibiting approximately 20% discrepancies for $\kappa\geq 0.2$, which falls well within the DM-dominated regime. Moreover, the lower the redshift, the smaller the $\kappa$ value at which this breakdown occurs. Interestingly, for the shear, we observe much better agreement between PINOCCHIO and $N$-body in the region of interest. Differences become apparent at $\gamma\gtrsim 0.1$, a region dominated by baryonic effects. An explanation for this disagreement will be proposed in Section 3.5. For the $\mu-$PDFs, we again observe that as we venture into more non-linear regimes, specifically for $\mu>2$, notable differences emerge between both methods. However, these differences are proportionally smaller in the DM-dominated regime than in the case of $\kappa$. Particularly, the agreement is consistently good for $\mu<1$—where the PDF rapidly decreases, causing relative residuals to increase—and ranges between $20\%$ and $40\%$ for larger $\mu$ until baryonic physics becomes dominant. As we will explore later, the multiple-image regime begins beyond these baryonic thresholds; thus, only the so-called type-I $\mu$-PDF is relevant here (see, e.g., Takahashi et al., 2011; Castro et al., 2018). A similar comparison for different angular resolutions is presented in Appendix A.
@@ 3.2 Strong lensing PDFs
Figure 4: Convergence, shear and magnification PDFs at $z=5$ for different angular resolutions for the different image types.
@@
Figure 4 displays the PDFs of $\kappa$, $\gamma$, and $\mu$ for $z=5$, categorized into these three types. We see that the $N$-body simulations show a higher probability of forming Types II and III images compared to structures generated by the PINOCCHIO code, highlighting the limitations of the method in accurately describing the strong-lensing regime (in the present implementation the halo internal structure is described by a single spherical NFW profile without substructure). The absence of most Type II and III images in PINOCCHIO is also consistent with observations from the previous section that the PINOCCHIO PDFs disagree from the $N$-body results for $\gamma\gtrsim 0.1$. Furthermore, we find that the angular resolution does not significantly impact the lensing PDFs in the strong-lensing regime, and comparisons between the PINOCCHIO and PINOCCHIOhalos PDFs show nearly indistinguishable results, with the exception of the Type II $\kappa$ PDFs.
@@ 3.3 Dissecting the PDF
Next, we show, breaking the sums of Eqs. (16) and (17) into individual elements, the contributions to the variance $\mu_{2}$ and skewness $\tilde{\mu}_{3}$ of halos with a given mass (Fig. 6) and in a given lensing plane redshift (Fig. 7), for sources at $z=1$ and $z=3$. It is also shown the average number of lensing-halo hits per light ray. In the case of Fig. 7, a value of 10 in a given $\Delta z$, for instance, means that the light ray hit on average 10 halos in that redshift bin (bin widths are constant in comoving distance), and similarly for Fig. 6.
@@
Figure 6: Contribution to the variance $\mu_{2}$ (top) and skewness $\tilde{\mu}_{3}=\mu_{3}/\sigma^{3}$ (bottom) of lensing halos with a given mass for sources at $z=1$ and $z=3$ using Eqs. (16)–(17). Also shown (point markers) is the average number of lensing-halo hits per light ray. See Section 3.3.
@@
From Fig. 6 (top panel) we can see that our simulations consider the relevant range of masses as the halos with mass $\approx 10^{13}h^{-1}M_{\odot}$ contribute to most of the variance, and halos with masses below $10^{9}h^{-1}M_{\odot}$ or above $10^{15}h^{-1}M_{\odot}$ have a negligible impact. Also, the lensing PDF for sources at $z=3$ has a variance which is about 5 times greater than the PDF for sources at $z=1$ – the light ray encounters more lenses in the former case. Regarding skewness, we observe that more massive halos contribute significantly to the overall asymmetry of the PDF. This is due to less massive halos being more numerous, thereby distributing the same fraction of the Universe’s mass across a greater number of objects. Consequently, a ray experiences a higher number of interactions (hits). According to the central limit theorem, this results in a PDF that more closely approximates a Gaussian distribution. Finally, a source at $z=3$ is further away, and a larger number of halos are encountered so that the number of hits is larger and the skewness is lower. At the very high-mass end, the skewness for sources at $z=3$ is slighter higher because the extra number of hits with respect to $z=1$ does not compensate for the very skewed distributions of those very massive halos.
@@
The top panel of Fig. 7 shows the contributions of the lens planes to the variance. The shape is modulated by the lensing efficiency of Eq. (7) and the growth of structures. The largest contributions come at about half way between source and observer. The bottom panel shows the contribution to the skewness. The overall trend is that lens planes at lower redshifts contribute to a higher skewness. From Fig. 6, we remember that smaller halos have a smaller number of hits at lower redshifts (and produce relatively more skewness), while this is the opposite for larger halos. The overall effect is dominated by the smaller halos, which can be attributed to the progressive nonlinearity in the Universe due to the growth of structures. We also see that the contributions for different source redshifts are exactly one on top of the other. It is useful to remember that the skewness in the matter and lensing distributions is due to the fact that under-densities occupy more volume than overdensities, so it is more likely to encounter an under-density or have a de-magnified light beam. Therefore, the third standardized moment is only due to geometry and does not depend on the lensing efficiency of Eq. (7), which depends on the source redshift.
@@
Figure 7: Contribution to the variance (top) and skewness (bottom) of halos of a given lensing plane redshift for sources at $z=1$ and $z=3$ using Eqs. (16)-(17). Also shown (point markers) is the average number of lensing-halo hits per light ray. See Section 3.3.
3.4 Momenta of the magnitude PDF
@@
We now focus on analyzing the second, third, and fourth central moments of the shift in magnitude $\Delta m$, which directly relates to magnification $\mu$ through:
$\Delta m\equiv 2.5\log_{10}\mu.$ (30)
@@
Using magnitudes is advantageous because the magnitude PDF exhibits more stable central moments than the $\mu$–PDF, which has a divergent variance in cosmology. Nevertheless, truncating the magnitude PDF at a certain negative $\Delta m$ value is prudent, as highly magnified objects could, in principle, be individually detected and excluded from the statistics either a posteriori or during target selection a priori. [^12] These methods accurately recover the mean of the non-truncated PDFs, which is essential for unbiased cosmological analyses. Since the mean is a non-central moment, we will not discuss here the effects of truncation on it. Furthermore, as previously demonstrated, the baryonic contribution becomes dominant for $\mu\gtrsim 3$, which is outside the focus of this study.
@@
Results for redshifts 1 and 3, computed in the source plane with a fixed angular resolution of 1.76 arcsec, are depicted in Fig. 8. These PDFs are assessed up to a cutoff magnitude, $\Delta m_{\rm cut}$, beyond which the PDF is considered zero. The second central moment shows the strongest agreement, indicating PINOCCHIO’s capability to accurately reproduce the body rather than the tail of the PDF distribution – a consistency noted in earlier sections. Particularly, at $z=3$, PINOCCHIO and PINOCCHIOhalos agree more closely with $N$-body results compared to $z=1$, except for $\mu_{2,\rm{lens}}$ in PINOCCHIOhalos. We interpret the latter disagreement as due to the higher relative importance of non-halo particles at higher redshifts, at which halos give a less complete description of the matter field (see next section). This indicates that non-halo particles have a significant impact on the central moments, suggesting that the discrepancies with the $N$-body results are not simply due to the halo profiles and positioning.
Overall, PINOCCHIO well models statistical quantities within limits where baryonic effects are negligible. This precision makes it a useful tool for constructing simple models of central moments as functions of redshift and cosmological parameters. These models are crucial for the MeMo method (Quartin et al., 2014), which utilizes the weak lensing effect on the intrinsic distribution of magnitudes to constrain key cosmological parameters, such as the amplitude of the matter power spectrum and the matter density parameter (Quartin et al., 2014; Castro and Quartin, 2014; Castro et al., 2016a).
@@
Figure 8: Central moments’ ratio of the magnitude PDFs of the approximative methods with respect to the $N$-body simulation, for different values of $\Delta m_{\rm cut}$ and for $z=1$ and $z=3$. The blue (green) shaded regions indicate the limit where the baryonic effects is dominant, above 50%, for $z=1$ ($z=3$).

### v2

@@ 2.6 The $\mathcal{L}_{1}$ norm
Figure 1: Convergence PDFs for the different lensing methods considered in this study, for an angular resolution of 1.76 arcsec. The Poisson errors in the PDFs were not shown because their value is smaller than the scale of the figure.
Figure 2: Comparison of $\kappa$, $\gamma$ and $\mu$ PDFs (in the source plane) for different redshifts for the $N$-body and PINOCCHIO case for the angular resolution of 1.76 arcsec. $\texttt{\small PINOCCHIO}_{\textrm{halos}}$ (not shown here for compactness) exhibits similar behavior to PINOCCHIO. We use semi-log scale for $\kappa$ and log-log for $\gamma$ and $\mu$ for better clarity. The shaded regions mark where baryonic corrections exceed 50% for $z=1$ (blue) or for $z=3$ (green). As can be seen, the two methods agree reasonably well in the range where baryonic effects are under control. See also Fig. 3.
@@ 3 Results
[…]
@@ 3.1 Lensing PDFs
However, differences in the tails of the distributions are obscured by the linear scale used in Fig. 1. Therefore, in Fig. 2, we present a general comparison between the results from $N$-body and PINOCCHIO for $\kappa$, $\gamma$, and $\mu$ on a logarithmic scale, depicting four different redshifts for each method. The turboGL code is excluded from this comparison as it does not compute the shear PDFs so that $\mu$ can be inferred only in the weak-lensing regime. As previously discussed, our analysis focuses on the region where baryonic effects are maintained below the 50% level. The regimes where this is not the case at $z=1$ and $z=3$ are highlighted with a blue and green shade, respectively. It is evident that as the source distance increases, the convergence, shear, and magnification PDFs become wider (indicating higher variance), and the peaks of their distributions shift (indicating skewness). We will further explore these trends for convergence in Section 3.3, noting the general trend that lensing intensifies at higher redshifts due to the increased number of lens encounters. We find that both PINOCCHIO and $N$-body simulations are consistent with the theoretical model of randomly distributed point-mass lenses (Schneider et al., 1992), which predicts that when $\gamma<0.01$, the PDF is proportional to $\gamma^{2}$, and for $\mu\gg 1$, the PDF is proportional to $\mu^{-1}$. However, while the asymptotic behavior prediction of the randomly distributed points model for the $\gamma-$PDF happens within the range where the baryonic effects are subdominant, the prediction for the $\mu-$PDF occurs in the regime deeply dominated by baryonic physics and is not shown in Fig. 2. The behavior of the magnification PDF at this regime will be discussed in Sect. 3.2. We observe a good agreement between PINOCCHIO and $N$-body for the shear, with differences becoming apparent at $\gamma\gtrsim 0.1$, a region dominated by baryonic effects. An explanation for this disagreement will be proposed in Section 3.5.
Figure 3 uses $\mathcal{L}_{1}(-\infty,x)$ to illustrate the differences between the PDFs from PINOCCHIO and turboGL compared to the $N$-body results. Regarding convergence, PINOCCHIO exhibits a small excess of probability density at the peak of the distribution, leading to a 3% relative error. turboGL shows similar agreement at $z=3$, which becomes worse at $z=1$. Regarding magnification, PINOCCHIO again shows a 3% error. The discrepancy on the shear PDF is around 2%, and roughly half of that if we limit $\gamma\lesssim 0.1$.
We show in Table 4 a comprehensive comparison of the various approximations. The first two lines summarize the results of Figure 3. The third line shows that the ${\cal{L}}_{1}$ norm is not sensitive to the tails of the distributions, where type 2 and 3 images are relevant (see next Section). The fourth line shows that non-collapsed matter plays a smaller but non-negligible role in the lensing PDFs, especially at higher redshifts (see also Section 3.5). The fifth line shows that the use of the Born approximation leads to further small deviations for the bulk of the PDFs. These are smaller changes than those introduced by PINOCCHIO, but if one is interested in recovering the full accuracy of an expensive Hydro simulation, we find that it may be appropriate to adopt a full, no Born method. Finally, the last line shows that, compared to the difference between $N$-body and PINOCCHIO, the difference between Hydro and $N$-body is similar for $\kappa$ and $\mu$, and three times as large for $\gamma$.
The larger difference between hydro and $N$-body simulations for $\gamma$ compared to $\kappa$ and $\mu$ observed in Table 4 is due to the contributions of multiple images events (discussed in the next section). For the shear PDF, these contributions cluster over a small dynamic range $10.1\lesssim\gamma<$ 1, while for the other variables, they are sparsed over a much more extensive range. Therefore, the crossing of the $50\%$ threshold is steeper for shear than for the other variables, causing a more significative impact in calculating $\mathcal{L}_{1}$.
Refer to caption Refer to caption Refer to caption
Figure 3: The $\mathcal{L}_{1}(-\infty,x)$ norm, where $x=\kappa,\,\gamma$ or $\mu$, computed between approximate methods and the full $N$-body results in the source plane, for $z=1$ (solid lines) and $z=3$ (dashed lines). The blue (green) shaded regions represent the interval in which the baryonic correction is over 50% for $z=1$ ($z=3$).
Table 4: Values of the ${\cal{L}}_{1}$ norm for the different cases.
Reference Surrogate $z=1\,[\%]$ $z=3\,[\%]$
$\kappa$ $\gamma$ $\mu$ $\kappa$ $\gamma$ $\mu$
$N$-body tgl 6.6 3.3
$N$-body PINOCCHIO 2.8 1.4 2.8 2.5 2.6 2.5
$N$-body PINOCCHIO 2.8 1.4 2.8 2.5 2.6 2.5
(only type I)
PINOCCHIO PINOCCHIOhalos 1.1 0.5 1.1 2.0 1.5 1.9
no Born Born 1.3 1.2 1.3 1.0 1.6 1.1
Hydro $N$-body 1.3 4.3 1.3 2.6 7.3 2.6
@@ 3.2 Strong lensing PDFs
Figure 4: Convergence, shear, and magnification PDFs for different image types at $z=5$ and an angular resolution of 1.76 arcsec.
@@
Figure 4 displays the PDFs of $\kappa$, $\gamma$, and $\mu$ for $z=5$, categorized into these three types. We see that the $N$-body simulations show a higher probability of forming Types II and III images compared to structures generated by the PINOCCHIO code, highlighting the limitations of the method in accurately describing the strong-lensing regime (in the present implementation, the halo internal structure is defined by a single spherical NFW profile without substructure). The absence of most Type II and III images in PINOCCHIO is also consistent with observations from the previous section that the PINOCCHIO PDFs disagree with the $N$-body results for $\gamma\gtrsim 0.1$. Furthermore, we find that the angular resolution does not significantly impact the lensing PDFs in the strong-lensing regime, and comparisons between the PINOCCHIO and PINOCCHIOhalos PDFs show nearly indistinguishable results, except the Type II $\kappa$ PDFs.
@@ 3.3 Dissecting the PDF
Next, we show, breaking the sums of Eqs. (16) and (17) into individual elements, the contributions to the variance $\mu_{2}$ and skewness $\tilde{\mu}_{3}$ of halos with a given mass (Fig. 6) and in a given lensing plane redshift (Fig. 7), for sources at $z=1$ and $z=3$. Here, we only consider the 1-halo term. It is also shown the average number of lensing-halo hits per light ray. In the case of Fig. 7, a value of 10 in a given $\Delta z$, for instance, means that the light ray hit on average 10 halos in that redshift bin (bin widths are constant in comoving distance), and similarly for Fig. 6.
@@
Figure 6: Contribution to the variance $\mu_{2}$ (top) and skewness $\tilde{\mu}_{3}=\mu_{3}/\sigma^{3}$ (bottom) of lensing halos with a given mass for sources at $z=1$ and $z=3$ using Eqs. (16-17). Also shown (point markers) is the average number of lensing-halo hits per light ray. See Section 3.3.
@@
From Fig. 6 (top panel) we can see that our simulations consider the relevant range of masses as the halos with mass $\approx 10^{13}h^{-1}M_{\odot}$ contribute to most of the variance. Halos with masses below $10^{9}h^{-1}M_{\odot}$ or above $10^{15}h^{-1}M_{\odot}$ have a negligible impact. This also means that the geometrical optics limit, which we assume here, should be valid for any source with wavelengths $\lesssim 10^{10}$(Leung et al., 2023), which covers not only supernovae but also future gravitational wave sources detected on the ground and on space with LISA. We also note that the lensing PDF for sources at $z=3$ has a variance that is about 5 times greater than the PDF for sources at $z=1$ – the light ray encounters more lenses in the former case. Regarding skewness, we observe that more massive halos contribute significantly to the overall asymmetry of the PDF. This is due to less massive halos being more numerous, thereby distributing the same fraction of the Universe’s mass across a greater number of objects. Consequently, a ray experiences a higher number of interactions (hits). According to the central limit theorem, this results in a PDF that more closely approximates a Gaussian distribution. Finally, a source at $z=3$ is further away, and a larger number of halos are encountered so that the number of hits is larger and the skewness is lower. Additionally, in this case, the lensing efficiency peaks at a higher redshift — $z\approx 1.5$ compared to $z\approx 0.5$ — making lensing more sensitive to structures that are more linear and on larger scales, thus closer to Gaussianity. At the very high-mass end, the skewness for sources at $z=3$ is slighter higher because the extra number of hits with respect to $z=1$ does not compensate for the very skewed distributions of those very massive halos.
@@
The top panel of Fig. 7 shows the contributions of the lens planes to the variance. The shape is modulated by the lensing efficiency of Eq. (7) and the growth of structures. The largest contributions come at about half way between source and observer. The bottom panel shows the contribution to the skewness. The overall trend is that lens planes at lower redshifts contribute to a higher skewness. From Fig. 6, we remember that smaller halos have a smaller number of hits at lower redshifts (and produce relatively more skewness), while this is the opposite for larger halos. The overall effect is dominated by the smaller halos, which can be attributed to the progressive nonlinearity in the Universe due to the growth of structures. We also see that the contributions for different source redshifts are exactly one on top of the other, suggesting that the third standardized moment is only due to geometry – underdensities occupy more volume than overdensities – and does not depend on the lensing efficiency. Indeed, from Eqs. (16)-(17), considering the contribution from the redshift bin $i$, we have that:
$\tilde{\mu}_{3}^{i}=$ $\frac{\sum_{u}\kappa_{1iu}^{3}\Delta N_{iu}}{\big{(}\sum_{u}\kappa_{1iu}^{2}\Delta N_{iu}\big{)}^{3/2}}=\frac{\sum_{u}\Sigma_{iu}^{3}\Delta N_{iu}}{\big{(}\sum_{u}\Sigma_{iu}^{2}\Delta N_{iu}\big{)}^{3/2}}\;,$ (32)
where, after using used Eq. (13), the lensing efficiency of Eq. (7) cancels out.
@@
Figure 7: Contribution to the variance (top) and skewness (bottom) of halos of a given lensing plane redshift for sources at $z=1$ and $z=3$ using Eqs. (16-17). Also shown (point markers) is the average number of lensing-halo hits per light ray. See Section 3.3.
3.4 Moments of the magnitude PDF
@@
We now focus on analyzing the second, third, and fourth central moments of the shift in magnitude $\Delta m$, which directly relates to magnification $\mu$ through
$\Delta m\equiv-2.5\log_{10}\mu.$ (33)
@@
Using magnitudes is advantageous because the magnitude PDF exhibits more stable central moments than the $\mu$–PDF, which has a divergent variance in cosmology. Nevertheless, truncating the magnitude PDF at a certain negative $\Delta m$ value is prudent, as highly magnified objects could, in principle, be individually detected and excluded from the statistics either a posteriori or during target selection a priori. [^14] These methods accurately recover the mean of the non-truncated PDFs, essential for unbiased cosmological analyses. Since the mean is a non-central moment, we will not discuss the effects of truncation here. Furthermore, as previously demonstrated, the baryonic contribution becomes dominant for $\mu\gtrsim 2.5$, which is outside the focus of this study. In other words, we use $\Delta m_{\rm cut}$ as a proxy for a series of systematic and sample-size effects on the observable PDF.
@@
Results for redshifts 1 and 3, computed in the source plane with a fixed angular resolution of 1.76 arcsec, are depicted in Fig. 8. These PDFs are assessed up to a cutoff magnitude, $\Delta m_{\rm cut}$, beyond which the PDF is considered zero. The second central moment shows the strongest agreement, indicating PINOCCHIO’s capability to accurately reproduce the body rather than the tail of the PDF distribution – a consistency noted in earlier sections. Particularly, at $z=3$, PINOCCHIO and PINOCCHIOhalos agree more closely with $N$-body results compared to $z=1$, except for $\mu_{2,\rm{lens}}$ in PINOCCHIOhalos. We interpret the latter disagreement as due to the higher relative importance of non-halo particles at higher redshifts, at which halos give a less complete description of the matter field (see next section). This indicates that non-halo particles significantly impact the central moments, suggesting that the discrepancies with the $N$-body results are not simply due to the halo profiles and positioning.
@@
Figure 8: Central moments’ ratio of the magnitude PDFs of the various methods with respect to the $N$-body simulation, for different values of $\Delta m_{\rm cut}$ and for $z=1$ and $z=3$. The blue (green) shaded regions indicate the limit where the baryonic effects are dominant, above 50%, for $z=1$ ($z=3$).
@@
Fig. 8 is also useful to quantify the impact of baryons in a way that is complementary to the ${\cal{L}}_{1}$ norm of Table 4. We see that for $\Delta m_{\rm cut}\simeq-0.4$, equivalent to $\mu\approx 1.5$, baryonic effects are negligible for the second moment and below 10% for the third and fourth moments. Furthermore, we can see that the 50% thresholds that we introduced in Table 3 correspond to 7%, 22% and 26% error for the second, third and fourth moments, respectively.
Overall, PINOCCHIO well models statistical quantities within limits where baryonic effects are negligible. This precision makes it useful for constructing simple models of central moments as functions of redshift and cosmological parameters. These models are crucial for the MeMo method (Quartin et al., 2014), which utilizes the weak lensing effect on the intrinsic distribution of magnitudes to constrain key cosmological parameters, such as the amplitude of the matter power spectrum and the matter density parameter (Quartin et al., 2014; Castro and Quartin, 2014; Castro et al., 2016b).